# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ Abstract
Quantum networks allow for novel forms of quantum nonlocality. By exploiting the combination of entangled states and entangled measurements, strong nonlocal correlations can be generated across the entire network. So far, all proofs of this effect are essentially restricted to the idealized case of pure entangled states and projective local measurements. Here we present noise-robust proofs of network quantum nonlocality, for a class of quantum distributions on the triangle network that are based on entangled states and entangled measurements. The key ingredient is a result of approximate rigidity for local distributions that satisfy the so-called “parity token counting” property with high probability. Considering quantum distributions obtained with imperfect sources, we obtain a noise robustness up to $\sim 80\%$ for dephasing noise and up to $\sim 0.5\%$ for white noise. Additionally, we can prove that all distributions in the vicinity of some ideal quantum distributions are nonlocal, with a bound on the total-variation distance. Our work opens interesting perspectives towards the practical implementation of quantum network nonlocality.
I Introduction
A growing interest has recently been devoted to the question of quantum nonlocality in networks, see e.g. [1] for a review. A general framework has been developed to investigate nonlocal correlations in networks featuring independent sources [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. A central motivation comes from the fact that the network setting allows for novel forms of nonlocality compared to the more standard Bell test, where a single source distributes a physical resource (e.g. an entangled quantum state) to the parties.
@@ I Introduction
Another key aspect of the network setting is the possibility to exploit joint entangled measurements (where one or more eigenstates are non-separable) [13]. Similarly, as in quantum teleportation, the combination of entangled states and entangled joint measurements allows for strong correlations to be distributed across the entire network. Notably, this leads to novel forms of quantum nonlocal correlations that are genuine to the network structure [14], in the sense that they crucially rely on the use of entanglement joint measurements [15, 16]. This is possible for a simple triangle network, as in Fig. 1.
@@
While these works represent significant progress in our understanding of quantum nonlocality, they suffer from one major limitation. Namely, these proofs of genuine quantum network nonlocality are derived in an ideal noiseless setting, where the sources distribute ideal (pure) quantum states and nodes perform ideal (projective) quantum measurements. However, these proofs no longer apply when any realistic noise level is considered. This represents a major hurdle towards any practical implementation or application of these ideas, as any experiment features a certain amount of noise originating from unavoidable technical imperfections. Moreover, numerical analysis based on neural networks indicates noise-robust nonlocality [17].
@@
One can draw a parallel with the situation of nonlocality in the mid-sixties. Indeed, John Bell’s initial proof of nonlocality applied only to the idealized case of a pure entangled state and projective local measurements [18]. A key step was then made by Clauser, Horne, Shimony and Holt who derived a noise-robust proof of quantum nonlocality via their famous CHSH Bell inequality [19], which enabled the first Bell experiments. This is still today the most commonly used test of nonlocality, and the basis of many applications within the device-independent approach (see e.g. [20]).
In this work we address the question of noise-robust proofs of quantum network nonlocality. We derive methods for this problem and discuss several examples. Our main contributions are the following. We present a general family of quantum distributions on the triangle network (without inputs) and prove their nonlocality in the ideal (noiseless) case. These distributions are based on the concept of “parity token counting” (PTC) and include as a special case the distributions of Renou et al. [14], from now on referred to as RGB4. In turn, we derive noise-robust nonlocality proofs for these quantum distributions. The key ingredient is a result of approximate rigidity for local distributions that are PTC, which allows us to deal with noise. In particular, we identify quantum distributions that remain nonlocal when up to $\sim 80\%$ dephasing noise, or $\sim 0.67$% white noise, is added at each source. Moreover, we characterize the distributions in the vicinity of a given nonlocal distribution. For a specific quantum distribution, we can show that any distribution in a $\sim 0.27$% total-variation distance ball around it is nonlocal. This can be viewed as a Bell inequality, in the sense that it guarantees that a full-measure region of the correlation set is nonlocal.
The paper is organized as follows. In Section II, we provide the general context, introduce notations, and review the concept of parity token counting (PTC) distributions. In Section III, we prove results on approximate PTC rigidity for local distributions. This allows us to bound local correlations that approximately verify the PTC condition. Additionally, we derive an analytical outer relaxation of the set of binary output correlations compatible with a triangle-local model. In Section IV, we introduce a general family of quantum nonlocal distributions. We first prove their nonlocality in the noiseless case. In turn, based on the results of Section III, we prove results on the nonlocality of these distributions in the noisy case. Finally, we conclude in Section VI and discuss future perspectives.
II Background
@@
To start we recall some background notions and fix the notations. Consider a triangle network as in Fig. 1, where three independent sources (labeled $\alpha$, $\beta$, and $\gamma$) distribute physical systems to three distant parties (nodes $A$, $B$, and $C$). Each party thus receives two physical subsystems (from two different sources), and produces a measurement outcome, denoted by the classical variables $a$, $b$, and $c$ (of finite cardinality $d$  [^1] For simplicity we assume here that the cardinality is the same for all outputs.). The experiment results in the joint probability distribution $P(a,b,c)$.
@@
where $\mathds{E}$ is the expected value over the independently distributed classical variables $\alpha,\beta,\gamma$. Furthermore, the conditional probabilities $P_{X}(x|\xi,\xi^{\prime})$ can be considered deterministic without loss of generality, and represented by the response functions $x(\xi,\xi^{\prime})$. A distribution that does not admit a decomposition of the form (1) is termed triangle-nonlocal, or simply nonlocal in the following.
@@
where $\rho^{\xi}$ denote the quantum states distributed by the sources (density operators), and $\{E_{YY^{\prime}}^{x}\}$ denote the measurements performed at each node (POVMs). Note that one should pay attention to the order of subsystems when computing (2).
@@
For $d\geq 3$, it is known that $\mathcal{L}_{\triangle}^{d}\subset\mathcal{Q}_{\triangle}^{d}$, i.e. there exists quantum distributions that are nonlocal (see [22] and references therein). For $d=2$, this is still an open question.
@@
Known proofs of quantum nonlocality for the triangle network are of two types. The first are distributions that can be viewed as a clever embedding of a standard Bell test (e.g. CHSH) on the triangle network without inputs. The initial construction was derived by Fritz [4]; see also [23, 24, 25, 22]. For this class of distributions, noise-robust methods have been recently developed [26], leading to a first experiment [12]. Due to their connection with standard Bell tests, these distributions can be implemented without the need for any entangled measurements. Hence the nonlocality of these distributions originates solely from the use of an entangled state.
@@
In contrast, there exist quantum distributions on the triangle network based on the judicious combination of entangled states and measurements, as first shown by Renou and colleagues [14]. It was recently shown that this distribution requires in fact the use of entangled measurements [16], hence demonstrating genuine network quantum nonlocality [15]. So far, these nonlocality proofs are essentially restricted to an ideal noiseless scenario see e.g. [14, 27, 28, 22, 29]). That is, they consider a setting where the shared states $\rho^{\xi}$ are pure and the local measurements $\{E_{YY^{\prime}}^{x}\}$ are projective. However, these proofs no longer apply when any realistic amount of noise is included [^2] Note the semi-analytical methods of inflation [6] can detect a specific instance of the RGB4 distribution, but the resulting noise-robustness is around $10^{-7}$ for white noise at the sources, so irrelevant from any practical perspective.. On the other hand, numerical analysis based on machine learning suggests that the nonlocality of these quantum distributions is in fact robust to noise [17], up to $\approx 10\%$ white noise at the sources for some instances of the RGB4 distribution. Deriving noise-robust analytical proofs is therefore an important open problem, which we address in this work.
II.1 PTC distributions and rigidity
@@
A key class of distributions for our work is termed parity token counting (PTC). These are local distributions on the triangle network, obtained via the following model. Each source possesses a single token that it sends to either one of its connected parties with some given probability. For instance, the source $\alpha$ sends its token to Bob with probability $p_{\alpha}$ and to Charlie with probability $1-p_{\alpha}$. Then each party outputs the parity of the total number of received tokens. Therefore if we define the binary random variable $t_{\alpha}\in\{0,1\}$ as the number of tokens $\alpha$ sends to Bob, $t_{\beta}\in\{0,1\}$ as the number of tokens $\beta$ sends to Charlie, and $t_{\gamma}\in\{0,1\}$ as the number of tokens $\gamma$ sends to Alice, then for example Alice is receiving $t_{\gamma}$ from the source $\gamma$ and $t_{\beta}\oplus 1$ from the source $\beta$ and its response function woud be $a(\beta,\gamma)=t_{\gamma}\oplus t_{\beta}\oplus 1$, where $\oplus$ denotes the sum modulo two.
@@
$\begin{split}&T_{\xi}:\xi\to t_{\xi}\in\{0,1\}\qquad\text{such that}\\ &x(\xi,\xi^{\prime})=t_{\xi}\oplus t_{\xi^{\prime}}\oplus 1.\end{split}$ (4)
In addition, if the probability distribution is such that $E_{a},E_{b},E_{b}\neq 0$ for $E_{x}=\text{Pr}(x=0)-\text{Pr}(x=1)$, the token functions can be chosen to fulfill
@@
and similar equalities of the other sources. This means that there is essentially only a unique local model over the triangle which can simulate such a distribution. In [22] one can find a simple characterization of the local PTC set $\mathcal{L}_{\triangle|PTC}^{2}$, containing all triangle-local distributions that satisfy the PTC condition of Eq. (3). We find that this set covers about 12.5% of the volume [^3] We have computed the volume of the set by uniformly sampling $10^{6}$ points $P$ from the PTC slice $\mathcal{P}_{PTC}^{2}$ and verifying for each point if it admits a local model or not. For symmetric distributions $E_{A}=E_{B}=E_{C}$ we find that the set $\mathcal{L}_{\triangle|PTC}^{2}$ is characterized by a simple inequality $P(1,1,1)\geq\frac{1}{4}$. of $\mathcal{P}_{PTC}^{2}$ – the set of binary probability distributions $P(a,b,c)$ fulfilling Eq. (3).
Remarkably the PTC rigidity, albeit formulated differently, is also true for any quantum PTC distribution [16]. In particular, this implies that in the slice of the probability set where the parity condition in Eq. (3) holds there is no separation between the quantum and local sets, i.e. $\mathcal{Q}_{\triangle|PTC}^{2}=\mathcal{L}_{\triangle|PTC}^{2}$.
III Approximate PTC rigidity for local distributions
@@
follows form the parity condition at $(\alpha,\beta,\gamma_{*})\in\Lambda$. In addition, we have a lower bound on the probability that $(\alpha,\beta)$ are sampled in $G_{*}$
@@
For any point $(\alpha,\beta,\gamma)$ from this set we can define the prolongation of the token function $t_{\gamma}$ as
@@
where the equality of the two definitions follows from the fact that both points $(\alpha,\beta,\gamma)$ and $(\alpha,\beta,\gamma_{*})$ are inside $\Lambda$ by construction. It is also easy to see that this value $t_{\gamma}$ satisfy the PTC rigidity, i.e. $t^{(\alpha,\beta)}_{\gamma}\oplus t_{\beta}\oplus 1=a(\beta,\gamma)$ and the other equations, for the variables value $(\alpha,\beta,\gamma)$ (see Appendix A). The problem however is that the prolongations of $t_{\gamma}$ must not necessarily agree for different choices of $(\alpha,\beta)$. Hence this strategy can not be used to define the token function $t_{\gamma}$ on the whole set $\Omega$. Nevertheless, we show in the Appendix A, that there exists a subset $\Upsilon\subset\Omega$ such that
$\text{Pr}(\bm{\xi}\in\Upsilon)$ $\geq 1-\text{f}_{\varepsilon}\geq 1-3\varepsilon\quad\text{with}$ (15)
$1-\text{f}_{\varepsilon}$ =$=\begin{cases}\frac{1}{2}-\varepsilon+\frac{1-4\varepsilon}{2\sqrt{1-2\varepsilon}}&\varepsilon\leq\frac{1}{4}\\ \frac{1-2\varepsilon}{2}&\varepsilon>\frac{1}{4}\end{cases}$ (16)
on which the token function $t_{\gamma}$ can be prolonged consistently. The key idea is to use the fact that the definitions in Eq. (14) must agree for any two points in $\Omega$ with the same $\alpha$ or $\beta$.
This demonstrates the following result holds.
[…]
The set $\Upsilon$ gives rise to the following subnormalized distribution $P(a,b,c,\bm{\xi}\in\Upsilon)$ which satisfies $\sum_{a,b,c}P(a,b,c,\bm{\xi}\in\Upsilon)\geq 1-\text{f}_{\varepsilon}$. For simplicity let us assume that this bound is tight (otherwise one can repeat the following discussion with a second parameter $\text{f}^{\prime}\leq\text{f}_{\varepsilon}$ and reach the same conclusions), it follows that
$P(a,b,c,\bm{\xi}\in\Upsilon)+\text{f}_{\varepsilon}P^{\prime}(a,b,c)=P(a,b,c),$ (17)
@@
Now, let us prolong the token functions $T_{\xi}$ to all the values of the local variables that do not appear in $\Upsilon$, this can be done by e.g. assigning $t_{\xi}=0$ to all such values. With the token functions defined everywhere, we can now define a new probability distribution
$\widetilde{P}(a,b,c)=\mathds{E}\Big{(}P_{\scriptscriptstyle ptc}(a|t_{\beta},t_{\gamma})P_{\scriptscriptstyle ptc}(b|t_{\alpha},t_{\gamma})P_{\scriptscriptstyle ptc}(c|t_{\alpha},t_{\beta})\Big{)}$ (18)
@@
By construction, this distribution is triangle local and PTC. Furthermore, we know that $P(a,b,c)$ and $\widetilde{P}(a,b,c)$ agree on the subset $\bm{\xi}\in\Upsilon$, hence the identity
$P(a,b,c,\bm{\xi}\in\Upsilon)+\text{f}_{\varepsilon}\widetilde{P}^{\prime}(a,b,c)=\widetilde{P}(a,b,c)$ (20)
holds for some PTC (non-necessarily local) distributions $\widetilde{P}^{\prime}$. Combining Eqs. (17,20) we conclude that the observed local distribution must satisfy $P(a,b,c)=\widetilde{P}(a,b,c)-\text{f}_{\varepsilon}\widetilde{P}^{\prime}(a,b,c)+\text{f}_{\varepsilon}P^{\prime}(a,b,c)$, which can be nicely expressed in terms of total variation distance

### v3

# These authors contributed equally to this work

@@ Abstract
Quantum networks allow for novel forms of quantum nonlocality. By exploiting the combination of entangled states and entangled measurements, strong nonlocal correlations can be generated across the entire network. So far, all proofs of this effect are essentially restricted to the idealized case of pure entangled states and projective local measurements. Here we present noise-robust proofs of network quantum nonlocality, for a class of quantum distributions on the triangle network that are based on entangled states and entangled measurements. The key ingredient is a result of approximate rigidity for local distributions that satisfy the so-called “parity token counting” property with high probability. Our methods can be applied to any type of noise. As illustrative examples, we consider quantum distributions obtained with imperfect sources and obtain a noise robustness up to $\sim 80\%$ for dephasing noise and up to $\sim 0.5\%$ for white noise. Additionally, we prove that all distributions in the vicinity of some ideal quantum distribution are nonlocal, with a bound on the total-variation distance $\sim 0.25\%$. Our work opens interesting perspectives towards the practical implementation of quantum network nonlocality.
1 Introduction
A growing interest has recently been devoted to the question of quantum nonlocality in networks, see e.g. [1] for a review. A general framework has been developed to investigate nonlocal correlations in networks featuring independent sources [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. A central motivation comes from the fact that the network setting allows for novel forms of nonlocality compared to the more standard Bell test, where a single source distributes a physical resource (e.g. an entangled quantum state) to the parties.
@@ 1 Introduction
Another key aspect of the network setting is the possibility to exploit joint entangled measurements (where one or more eigenstates are non-separable) [15]. Similarly, as in quantum teleportation, the combination of entangled states and entangled joint measurements allows for strong correlations to be distributed across the entire network. Notably, this leads to novel forms of quantum nonlocal correlations that are genuine to the network structure [16], in the sense that they crucially rely on the use of entangled joint measurements [17, 18]. This is possible for a simple triangle network, as in Fig. 1.
@@
While these works represent significant progress in our understanding of quantum nonlocality, they suffer from one major limitation. Namely, these proofs of genuine quantum network nonlocality are derived in an ideal noiseless setting, where the sources distribute ideal (pure) quantum states and nodes perform ideal (projective) quantum measurements. Formally, these proofs are based on the concept of “token counting” rigidity [16, 19, 20], which so far only applies to the ideal (noiseless) case. Hence, these nonlocality proofs no longer apply when any realistic noise level is considered. This represents a major hurdle towards any practical implementation or application of these ideas, as any experiment features a certain amount of noise originating from unavoidable technical imperfections. Moreover, numerical analysis based on neural networks indicates noise-robust nonlocality [21].
@@
One can draw a parallel with the situation of nonlocality in the mid-sixties. Indeed, John Bell’s initial proof of nonlocality applied only to the idealized case of a pure entangled state and projective local measurements [22]. A key step was then made by Clauser, Horne, Shimony, and Holt who derived a noise-robust proof of quantum nonlocality via their famous CHSH Bell inequality [23], which enabled the first Bell experiments [24, 25]. This is still today the most commonly used test of nonlocality [26, 27, 28], and the basis of many applications within the device-independent approach (see e.g. [29]).
In this work we address the question of noise-robust proofs of quantum network nonlocality. We derive methods for this problem and apply them to several examples of quantum nonlocal distributions that are based on the concept of “parity token counting” (PTC) [20]. Our main contribution is a result of approximate rigidity for local distributions that satisfy the PTC property with a high (but non-unit) probability. Importantly, this method can be applied to any type of noise. We present illustrative examples, considering quantum distributions obtained from imperfect sources, featuring either dephasing or white noise. Additionally, we show that all distributions in the vicinity of some ideal quantum distribution are nonlocal, with a bound on the total-variation distance, which can be viewed as a nonlinear Bell inequality for the triangle network. Our work opens interesting perspectives towards the practical implementation of Bell nonlocality in quantum networks.
2 Summary of results
To start, we first provide a summary of the main contributions of this work.
  1. 1.
    We present a general family of quantum distributions on the triangle network (without inputs) and prove their nonlocality in the ideal (noiseless) case, see Section 5. These distributions include as a special case the distributions of Renou et al. [16], from now on referred to as RGB4. The nonlocality proof is based on the “parity token counting” (PTC) rigidity property [20]. In words, it states that on the triangle network with binary outcomes, any local distribution for which the sum of the outputs is always odd (1 or 3) has a unique corresponding local model.
  2. 2.
    We derive noise-robust nonlocality proofs for these quantum distributions, see Section 5.2. Our methods can be applied to arbitrary type of noise. As an illustration, we demonstrate the nonlocality of quantum distributions where up to $\sim 80\%$ dephasing noise, or $\sim 0.55$% white noise, is added at each source. We also investigated other noise models, such as no-click and photon loss; see Appendix E. Moreover, in Section 5.3 we characterize the distributions in the vicinity of a given nonlocal distribution. For a specific quantum distribution, we can show that any distribution in a $\sim 0.25$% total-variation distance ball around it is nonlocal. This can be viewed as a Bell inequality, in the sense that it guarantees that a full-measure region of the correlation set (the ball around the quantum distribution) is nonlocal.
  3. 3.
    The key ingredient which allows us to prove noise-robust nonlocality, is a result on the approximate PTC rigidity, discussed in Section 4. It gives a quantitative characterization of all local models leading to distributions for which the PTC condition holds approximately. That is, the distributions for which the sum of the outputs is odd with high probability. With the help of this result, we also derive an analytical outer relaxation of the set of binary output correlations compatible with a triangle-local model.
Below we start with Section 3 where we introduce more formally the context, the notations, and review the concept of parity token counting (PTC) distributions.
3 Background
@@
To start we recall some background notions and fix the notations. Consider a triangle network as in Fig. 1, where three independent sources (labeled $\alpha$, $\beta$, and $\gamma$) distribute physical systems to three distant parties (nodes $A$, $B$, and $C$). Each party thus receives two physical subsystems (from two different sources), and produces a measurement outcome, denoted by the classical variables $a$, $b$, and $c$ (of finite cardinality $d$). The experiment results in the joint probability distribution $P(a,b,c)$.
@@
where $\mathds{E}$ is the expected value over the independently distributed classical variables $\alpha,\beta,\gamma$ [30]. Furthermore, the conditional probabilities $P_{X}(x|\xi,\xi^{\prime})$ can be considered deterministic without loss of generality, and represented by the response functions $x(\xi,\xi^{\prime})$. A distribution that does not admit a decomposition of the form (1) is termed triangle-nonlocal. For brevity, we refer to triangle-(non)local correlations simply as (non)local throughout the text.
@@
where $\rho^{\xi}$ denote the quantum states distributed by the sources (density operators), and $\{E_{YY^{\prime}}^{x}\}$ denote the measurements performed at each party and are given by positive operator valued measures (POVMs). Note that one should pay attention to the order of subsystems when computing (2).
@@
For $d\geq 3$, it is known that $\mathcal{L}_{\triangle}^{d}\subset\mathcal{Q}_{\triangle}^{d}$, i.e. there exist quantum distributions that are nonlocal (see [20] and references therein). For $d=2$, this is still an open question.
@@
Known proofs of quantum nonlocality for the triangle network are of two types. The first are distributions that can be viewed as a clever embedding of a standard Bell test (e.g. CHSH) on the triangle network without inputs [4]. The idea is to distribute effective inputs via two classical sources. Each party will broadcast the received inputs via their output. The third source distributes the entangled state as in the usual Bell test. See also [32, 33, 34, 20] for more examples. For this class of distributions, noise-robust methods have been recently developed [35], leading to a first experiment [14]. Due to their connection with standard Bell tests, these distributions can be implemented without the need for any entangled measurements. Hence, the nonlocality of these distributions originates solely from the use of an entangled state.
@@
In contrast, there exist quantum distributions on the triangle network based on the judicious combination of entangled states and measurements, as first shown by Renou and colleagues [16]. It was recently shown that this distribution requires in fact the use of entangled measurements [18], hence demonstrating genuine network quantum nonlocality [17]. So far, these nonlocality proofs are essentially restricted to an ideal noiseless scenario; see e.g. [16, 19, 36, 20, 37]. That is, they consider a setting where the shared states $\rho^{\xi}$ are pure and the local measurements $\{E_{YY^{\prime}}^{x}\}$ are projective. However, these proofs no longer apply when any realistic amount of noise is included [^2] Note the semi-analytical methods of inflation [6] can detect a specific instance of the RGB4 distribution, but the resulting noise-robustness is around $10^{-7}$ for white noise at the sources, so irrelevant from any practical perspective.. On the other hand, numerical analysis based on machine learning suggests that the nonlocality of these quantum distributions is in fact robust to noise [21], up to $\approx 10\%$ white noise at the sources for some instances of the RGB4 distribution. Deriving noise-robust analytical proofs is therefore an important open problem, which we address in this work.
3.1 PTC distributions and rigidity
@@
A key class of distributions for our work is termed parity token counting (PTC). These are local distributions on the triangle network, obtained via the following model. Each source possesses a single token that it sends to either one of its connected parties with some given probability. For instance, the source $\alpha$ sends its token to Bob with probability $p_{\alpha}$ and to Charlie with probability $1-p_{\alpha}$. Then each party outputs the parity of the total number of received tokens. Therefore, consider that we define the binary random variable $t_{\alpha}\in\{0,1\}$ as the number of tokens $\alpha$ sends to Bob, $t_{\beta}\in\{0,1\}$ as the number of tokens $\beta$ sends to Charlie, and $t_{\gamma}\in\{0,1\}$ as the number of tokens $\gamma$ sends to Alice. Then for example, Alice receives $t_{\gamma}$ from the source $\gamma$ and $t_{\beta}\oplus 1$ from the source $\beta$ and her response function would be $a(\beta,\gamma)=t_{\gamma}\oplus t_{\beta}\oplus 1$, where $\oplus$ denotes the sum modulo two.
@@
$\begin{split}&T_{\xi}:\xi\to t_{\xi}\in\{0,1\}\qquad\text{such that}\\ &x(\xi,\xi^{\prime})=t_{\xi}\oplus t_{\xi^{\prime}}\oplus 1,\end{split}$ (4)
for the sources $\xi,\xi^{\prime}\in\{\alpha,\beta,\gamma\}$ and $x\in\{a,b,c\}$ representing the party connected to the sources $\xi,\xi^{\prime}$. In addition, if the probability distribution is such that $E_{a},E_{b},E_{c}\neq 0$ for $E_{x}=\text{Pr}(x=0)-\text{Pr}(x=1)$, the token functions can be chosen to fulfill
@@
and similar equalities of the other sources. This means that there is essentially only a unique local model over the triangle which can simulate such a distribution. In [20] one can find a simple characterization of the local PTC set $\mathcal{L}_{\triangle|PTC}^{2}$, containing all triangle-local distributions that satisfy the PTC condition of Eq. (3). A graphical representation of $\mathcal{L}_{\triangle|PTC}^{2}$ is given in Fig. 2, we find that it covers only about 12.5% of the volume [^3] We have computed the volume of the set by uniformly sampling $10^{6}$ points $P$ from the PTC slice $\mathcal{P}_{PTC}^{2}$ and verifying for each point if it admits a local model or not. For symmetric distributions $E_{A}=E_{B}=E_{C}$ we find that the set $\mathcal{L}_{\triangle|PTC}^{2}$ is characterized by a simple inequality $P(1,1,1)\geq\frac{1}{4}$. of $\mathcal{P}_{PTC}^{2}$ – the set of binary probability distributions $P(a,b,c)$ fulfilling Eq. (3).
Figure 2: A layer-by-layer representation of the set $\mathcal{L}_{\triangle|PTC}^{2}=\mathcal{Q}_{\triangle|PTC}^{2}$, containing all triangle-local (or equivalently quantum) distributions $P(a,b,c)$ with binary outcomes, satisfying the deterministic PTC constraint $\text{Pr}(a\oplus b\oplus c=1)=1$. Each layer depicts the set for a given value of $P(a=b=c=1)$. The distributions inside the set can be represented by the vector $\bm{v}=\big{(}P(1,0,0),P(0,1,0),P(0,0,1)\big{)}$, with the remaining nonzero probability $P(1,1,1)$ given by normalization. Each plot in the figure is a slice of the set for a fixed value of $P(1,1,1)$, i.e. it is a cut of the set with the plane define by the equation $\bm{v}\cdot(1,1,1)=1-P(1,1,1)$. In each layer, the outer triangle delimits the set of valid probabilities. A 3D representation of the set $\mathcal{L}_{\triangle|PTC}^{2}$ in printable format can be obtained from the authors upon request.
Remarkably, the PTC rigidity, albeit formulated differently, is also true for any quantum PTC distribution [18]. In particular, this implies that in the slice of the probability set where the parity condition in Eq. (3) holds (ideal case without noise), there is no separation between the quantum and local sets, i.e. $\mathcal{Q}_{\triangle|PTC}^{2}=\mathcal{L}_{\triangle|PTC}^{2}$.
4 Approximate PTC rigidity for local distributions
@@
follows from the parity condition at $(\alpha,\beta,\gamma_{*})\in\Lambda$. In addition, we have a lower bound on the probability that $(\alpha,\beta)$s are sampled in $G_{*}$
@@
For any point $(\alpha,\beta,\gamma)$ from this set, we can define the prolongation of the token function $t_{\gamma}$ as
@@
where the equality of the two definitions follows from the fact that both points $(\alpha,\beta,\gamma)$ and $(\alpha,\beta,\gamma_{*})$ are inside $\Lambda$ by construction. It is also easy to see that this value $t_{\gamma}$ satisfies the PTC rigidity, i.e. $t^{(\alpha,\beta)}_{\gamma}\oplus t_{\beta}\oplus 1=a(\beta,\gamma)$ and the other equations, for the variables value $(\alpha,\beta,\gamma)$ (see Appendix A). The problem however is that the prolongations of $t_{\gamma}$ must not necessarily agree for different choices of $(\alpha,\beta)$. Hence this strategy can not be used to define the token function $t_{\gamma}$ on the whole set $\Omega$. Nevertheless, we show in Appendix A, that there exists a subset $\Upsilon\subset\Omega$ such that
$\begin{split}\text{Pr}(\bm{\xi}\in\Upsilon)&\geq 1-\text{f}_{\varepsilon}\geq 1-3\varepsilon\quad\text{with}\\ 1-\text{f}_{\varepsilon}&=\begin{cases}\frac{1}{2}-\varepsilon+\frac{1-4\varepsilon}{2\sqrt{1-2\varepsilon}}&\varepsilon\leq\frac{1}{4}\\ \frac{1-2\varepsilon}{2}&\varepsilon>\frac{1}{4}\end{cases}\end{split}$ (15)
on which the token function $t_{\gamma}$ can be prolonged consistently. The key idea is to use the fact that the definitions in Eq. (14) must agree for any two points in $\Omega$ with the same $\alpha$ or $\beta$. This demonstrates the following result holds.
Result 1. (Approximate PTC rigidity) For any triangle-local model leading to a distribution $P(a,b,c)\in\mathcal{L}_{\triangle}^{2}$ that fulfills the approximate parity condition $\text{Pr}(a\oplus b\oplus c=1)\geq 1-\varepsilon$, there exists a subset $\Upsilon$ of local variable values $\bm{\xi}=(\alpha,\beta,\gamma)$ with $Pr(\bm{\xi}\in\Upsilon)\geq 1-\text{f}_{\varepsilon}$ (see Eq. 15), on which one can define the token functions $T_{\alpha},T_{\beta},T_{\gamma}$ that satisfy the PTC rigidity $a(\beta,\gamma)=T_{\beta}\oplus T_{\gamma}\oplus 1$ etc.
For any triangle-local model, the set $\Upsilon$ gives rise to the subnormalized distribution $P(a,b,c;\bm{\xi}\in\Upsilon)$ representing the probability of the event $\xi=(\alpha,\beta,\gamma)\in\Upsilon$ and $a(\beta,\gamma)=a,b(\alpha,\gamma)=b,c(\alpha,\beta)=c$” for the output values $a,b,c$. These probabilities satisfy $\sum_{a,b,c}P(a,b,c;\bm{\xi}\in\Upsilon)\geq 1-\text{f}_{\varepsilon}$. For simplicity let us assume that this bound is tight (otherwise one can repeat the following discussion with a second parameter $\text{f}^{\prime}\leq\text{f}_{\varepsilon}$ and reach the same conclusions). It follows that
$P(a,b,c;\bm{\xi}\in\Upsilon)+\text{f}_{\varepsilon}P^{\prime}(a,b,c)=P(a,b,c),$ (16)
@@
Now, let us extend the token functions $T_{\xi}$ to all the values of the local variables that do not appear in $\Upsilon$, this can be done by e.g. assigning $t_{\xi}=0$ to all such values. With the token functions defined everywhere, we can now define the corresponding extended probability distribution as follows
$\widetilde{P}(a,b,c)$ (17)
$=\mathds{E}\Big{(}P_{\scriptscriptstyle ptc}(a|t_{\beta},t_{\gamma})P_{\scriptscriptstyle ptc}(b|t_{\alpha},t_{\gamma})P_{\scriptscriptstyle ptc}(c|t_{\alpha},t_{\beta})\Big{)}$
@@
By construction, this distribution is triangle-local and PTC. Furthermore, we know that $P(a,b,c)$ and $\widetilde{P}(a,b,c)$ agree on the subset $\bm{\xi}\in\Upsilon$, hence the identity
$P(a,b,c;\bm{\xi}\in\Upsilon)+\text{f}_{\varepsilon}\widetilde{P}^{\prime}(a,b,c)=\widetilde{P}(a,b,c)$ (19)
holds for some PTC (non-necessarily local) distributions $\widetilde{P}^{\prime}$. Combining Eqs. (16,19) we conclude that the observed local distribution must satisfy $P(a,b,c)=\widetilde{P}(a,b,c)-\text{f}_{\varepsilon}\widetilde{P}^{\prime}(a,b,c)+\text{f}_{\varepsilon}P^{\prime}(a,b,c)$, which can be nicely expressed in terms of total variation distance